\vskip 8mm
\begin{thebibliography}{20}

\bibitem{DiVincenzo} See, for example, D.P. DiVincenzo,
{\em Science} {\bf 270}, 255 (1995).

\bibitem{teleport} C.H. Bennett, G. Brassard, C. Cr\'epeau,
       R. Jozsa, A. Peres,
       and W.K. Wootters, {\em Phys. Rev. Lett.} {\bf 70}, 1895
       (1993).\\
S. Albeverio and S.M. Fei,
          {\em Phys. Lett.} {\bf A 276}(2000)8-11.\\
G.M. D'Ariano, P.Lo Presti, M.F. Sacchi, {\it Phys. Lett.} {\bf A
272} (2000), 32.\\
S. Albeverio and S.M. Fei and W.L. Yang, {\em Commun. Theor.
Phys.} {\bf 38} 301-304 (2002); {\em Phys. Rev.} {\bf A 66} 012301
(2002).

\bibitem{dense} C.H. Bennett and S.J. Wiesner,
        {\em Phys. Rev. Lett.} {\bf 69}, 2881 (1992).

\bibitem{crypto}
A.~Ekert, {\em Phys. Rev. Lett.} {\bf 67}, 661 (1991).\\
D.~Deutsch, A.~Ekert, P. Rozsa, C. Macchiavello,
S. Popescu and A. Sanpera, {\em Phys. Rev. Lett.} {\bf 77}, 2818
(1996).\\
C.A. Fuchs, N. Gisin,
         R.B. Griffiths, C-S. Niu, and
         A. Peres, {\em Phys. Rev. A}, {\bf 56}, 1163 (1997).

\bibitem{568} V. Vedral, M. B. Plenio, M. A. Rippin and
P. L. Knight: Phys. Rev. Lett. 78, 2275 (1997).\\
V. Vedral and M. Plenio: Phys. Rev. A 57, 1619 (1998).\\
V.~Vedral, M.B.~Plenio, K.~Jacobs, and P.L.~Knight, {\em Phys.
Rev. A} {\bf 56}, 4452 (1997).

\bibitem{7}
C.H. Bennett, H.J. Bernstein, S. Popescu,
and B. Schumacher, {\em Phys.~Rev.~A} {\bf 53}, 2046 (1996).

\bibitem{Peres96a}
A. Peres, Phys. Rev. Lett. {\bf 76} 1413 (1996).\\
K. \.Zyczkowski and P. Horodecki, Phys. Rev. A {\bf 58},  883
(1998).

\bibitem{sw}
B. Schumacher and M.D. Westmoreland, {\it Relative entropy in
quantum information theory}, quant-ph/0004045.\\
K. Audenaert, B. De Moor, K.G. H. Vollbrecht and R.F. Werner
Phys. Rev. A 66, 032310 (2002).\\
K. Audenaert, J. Eisert, E. Jane, M.B. Plenio, S. Virmani and B. De Moor,
Phys. Rev. Lett. 87 (2001) 217902.

\bibitem{17} C. H. Bennett, D. P. DiVincenzo, J. Smolin, and
W. K. Wootters: Phys. Rev. A 54, 3824(1996).

\bibitem{HillWootters} S.~Hill and W.K.~Wootters, {\em Phys. Rev. Lett.}
{\bf 78}, 5022 (1997).\\
W.K.~Wootters, {\em Phys. Rev. Lett.} {\bf 80}, 2245 (1998).

\bibitem{th}
B.M. Terhal, and K.G.H. Vollbrecht, {Phys. Rev. Lett.} {\bf 85},
2625 (2000).

\bibitem{ent3}
S.M. Fei, J. Jost, X.Q. Li-Jost and G.F. Wang,
Phys. Lett. A 310(2003)333-338.

\bibitem{peresbook} A. Peres, ``Quantum Theory: Concepts and Methods'',
Kluwer Academic Publishers (1995).

\bibitem{10} R. F. Werner: Phys. Rev. A 40, 4277 (1989).\\
S. Popescu: Phys. Rev. Lett. 72, 797 (1994).\\
S. Popescu: Phys. Rev. Lett. 74, 2619 (1995).\\
M. Lewenstein, D. Bruss, J.I. Cirac, B. Kraus, M. Kus, J.
Samsonowicz, A. Sanpera and R. Tarrach, J. Mod. Opt. 47 (2000)
2841.\\
B.M. Terhal, Theor. Comput. Sci. 287 (2002) 313.\\
M. Horodecki, P. Horodecki and R. Horodecki, Springer
Tracts in Mod. Phy. 173 (2001) 151.

\bibitem{Bell64} J.S. Bell, Physics (N.Y.) 1 (1964) 195.

\bibitem{3hPLA223} M. Horodecki, P. Horodecki and R. Horodecki, Phys. Lett.
A 223 (1996) 1.

\bibitem{2hPRA99} M. Horodecki and P. Horodecki, Phys. Rev. A 59 (1999) 4206.

\bibitem{cag99} N.J. Cerf, C. Adami and R.M. Gingrich, Phys. Rev. A 60
(1999) 898.

\bibitem{nielson01} M.A. Nielsen and J. Kempe, Phys. Rev. Lett. 86 (2001)
5184.

\bibitem{ter00} B. Terhal, Phys. Lett. A 271 (2000) 319.\\
M. Lewenstein, B. Kraus, J.I. Cirac and P. Horodecki, Phys.
Rev. A 62 (2000) 052310.

\bibitem{dps} A.C. Doherty, P.A. Parrilo and F.M. Spedalieri, Phys. Rev.
Lett. 88 (2002) 187904.

\bibitem{ru02} O. Rudolph, {\it Further results on the cross
norm criterion for separability}, quant-ph/0202121.\\
K. Chen and L.A. Wu, Quant. Inform. Comp. 3 (2003) 193-202.

\bibitem{chenPLA02} K. Chen and L.A. Wu, Phys. Lett. A 306 (2002) 14.

\bibitem{grc}
S. Albeverio, K. Chen and S.M. Fei, {\it Generalized reduction
criterion for separability of quantum states}, SFB-preprint, 2002.

\bibitem{hlpre} P. Horodecki, M. Lewenstein, G. Vidal and I. Cirac, Phys.
Rev. A 62 (2000) 032310.\\
P. Horodecki, J.A. Smolin, B.M. Terhal and A.V. Thapliyal,
Theor. Comp. Sci. 292 (2003)589-596.\\
S. Albeverio, S.M. Fei and D. Goswami, {\em Phys. Lett.} {\bf A},
91-96(2001).\\
S.M. Fei, X.H. Gao, X.H. Wang, Z.X. Wang and K. Wu, {\em Phys.
Lett. A} {\bf 300} (2002)559-566.\\
S.M. Fei, X.H. Gao, X.H. Wang, Z.X. Wang and K. Wu, {\sf Int. J.
Quant. Inform.}, {\bf 1}(2003)37-49.

\bibitem{22n}
S. Karnas and  M. Lewenstein, Phys. Rev. A 64, 042313 (2001).\\
S.M. Fei, X.H. Gao, X.H. Wang, Z.X. Wang and K. Wu,
{\sf Phys. Rev.} {\bf A 68}(2003)022315.

\bibitem{222n}
S.M. Fei, X.H. Gao, X.H. Wang, Z.X. Wang and K. Wu, {\it Canonical
Form and Separability of PPT States in  ${\cal C}^2 \otimes {\cal
C}^2 \otimes {\cal C}^2 \otimes {\cal C}^N$ composite quantum
systems}, to appear in {\sf Commun. Theor. Phys.} (2003).


\end{thebibliography}

\def\ra{\rangle}
\def\la{\langle}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\ba{\begin{array}}
\def\ea{\end{array}}

\def\ra{\rangle}
\def\la{\langle}

\documentstyle[11pt]{article}
\input amssym.def
\begin{document}
\baselineskip=18pt \setcounter{page}{1}
\begin{center}
{\LARGE \bf Canonical Form and  Separability of PPT States in
${\cal C}^2 \otimes {\cal C}^M \otimes {\cal C}^N$ Composite
Quantum Systems}
\end{center}
\vskip 2mm

\begin{center}
{\normalsize Xiao-Hong Wang$^1$, Shao-Ming Fei$^{1,\  2}$,
  Zhi-Xi Wang$^1$, and Ke Wu$^1$  }
\medskip

\begin{minipage}{4.8in}
{\small \sl $ ^1$
Department of Mathematics, Capital  Normal University, Beijing,
China.}\\
{\small \sl $ ^2$ Institute of Applied Mathematics, University of
Bonn,  53115 Bonn, Germany}
\end{minipage}
\end{center}

\vskip 2mm
\begin{center}

\begin{minipage}{4.8in}

\centerline{\bf Abstract}
\bigskip

We investigate the canonical forms of positive partial
transposition (PPT) density matrices in ${\cal C}^2 \otimes {\cal
C}^M \otimes {\cal C}^N$ composite quantum systems with rank $N$.
A general expression for these PPT states are explicitly obtained.
From this canonical form a sufficient separability condition is
presented.

\vskip 9mm
Key words: Separability, Quantum entanglement
\vskip 1mm
PACS number(s): 03.67.Hk, 03.65.Ta, 89.70.+c
\end{minipage}
\end{center}

\vskip0.4cm

Quantum entangled states have become one of the key resources in
the rapidly expanding field of quantum information processing and
computation \cite{DiVincenzo,teleport,dense,crypto}. Nevertheless,
the study of physical character and mathematical structure of the quantum
entanglement is far from being satisfied. One even does not
have a general criterion to judge if a quantum (mixed) state is
entangled or not. For bipartite states a number of entanglement measures such
as entanglement of formation and distillation \cite{568,7}, negativity
\cite{Peres96a}, von Neumann entropy and relative
entropy \cite{568,sw} have been proposed.
However most proposed measures of entanglement involve
extremizations which are difficult to handle analytically. For
instance, explicit analytic formulae for
entanglement of formation \cite{17} have been found only for
a pair of qubits system \cite{HillWootters}, and for some symmetric
states \cite{th} and a class of special states \cite{ent3}.
For multipartite systems there is no well
defined measure of entanglement yet.

The separability problem for pure states is quite well
understood \cite{peresbook}.
Nevertheless, in real conditions, due to the interactions with
environment, one encounters mixed states rather than pure ones.
The manifestations of mixed-state entanglement can be very subtle \cite{10}.
The Bell inequalities satisfied by a
separable system give the first necessary condition for
separability \cite{Bell64}. Afterwards the Peres criterion
\cite{Peres96a} says that partial transpositions with respect to
one or more subsystems of a separable state $\rho$ are positive.
This criterion
was further shown to be also sufficient for bipartite systems in
${\cal C}^2\otimes {\cal C}^2$ and ${\cal C}^2\otimes {\cal C}^3$
\cite{3hPLA223}. The reduction criterion proposed independently in
\cite{2hPRA99} and \cite{cag99} gives another necessary criterion
which is equivalent to the Peres criterion for ${\cal C}^2\otimes
{\cal C}^N$ composite systems but is generally weaker. There are
many other necessary criteria such as majorization
\cite{nielson01}, entanglement witnesses
\cite{3hPLA223,ter00}, extension of Peres criterion \cite{dps}, matrix
realignment \cite{ru02}, generalized partial transposition
criterion (GPT) \cite{chenPLA02}, generalized reduced criterion
\cite{grc}. For low rank density matrices there are also some
necessary and sufficient criteria of separability
\cite{hlpre}.

The separability and entanglement in ${\cal C}^2\otimes {\cal
C}^2\otimes {\cal C}^N$ and ${\cal C}^2\otimes {\cal C}^3\otimes
{\cal C}^N$ composite quantum systems have been studied in terms
of matrix analysis on tensor spaces \cite{22n}. It is shown that
all such quantum states $\rho$ with positive partial
transpositions and rank $r(\rho)\leq N$ are separable. The
canonical form and a sufficient separable condition PPT states in
${\cal C}^2 \otimes {\cal C}^2\otimes {\cal C}^2 \otimes {\cal
C}^N$ with rank $N$ are given in \cite{222n}. In this article we
extend the results in \cite{22n} to the case of composite quantum
systems in ${\cal C}^2 \otimes {\cal C}^M\otimes {\cal C}^N$. We
give a canonical form of positive partial transposition (PPT)
states in ${\cal C}^2 \otimes {\cal C}^M\otimes {\cal C}^N$ with
rank $N$ and present a sufficient separability criterion.

A separable state in ${\cal C}^2_{\sc a}  \otimes {\cal C}^M_{\sc
b} \otimes {\cal C}^N_{\sc c}$ is of the form:
\begin{equation}
\rho_{\sc abc}=\sum_{i}p_{i}\rho^{i}_{\sc a}\otimes \rho^{i}_{\sc
b} \otimes \rho^{i}_{\sc c}, \label{sep}
\end{equation}
where $\sum_{i}p_{i}=1$, $0< p_i\leq 1$, $\rho^{i}_{\alpha}$ are
density matrices associated with the subsystems $\alpha$,
$\alpha={\sc a,b,c}$. In the following we denote by $R(\rho)$,
$K(\rho)$, $r(\rho)$ and $k(\rho)$ the range, kernel, rank,
dimension of the kernel of $\rho$, respectively.

We first derive a canonical form of PPT states in ${\cal C}^2_{\sc
a} \otimes {\cal C}^4_{\sc b} \otimes {\cal C}^N_{\sc c}$ with
rank $N$, which allows for an explicit decomposition of a given
state in terms of convex sum of projectors on product vectors. Let
$|0_A\ra$, $|1_A\ra$; $|0_B\ra$, $|1_B\ra$, $|2_B\ra$, $|3_B\ra$
and $|0_C\ra\,, \cdots \,,|N-1_C\ra$ be some local bases of the
sub-systems ${\sc a}$, ${\sc b}$ and ${\sc c}$ respectively.

{\bf Lemma 1. }\ \  Every PPT state $\rho$ in ${\cal C}^2_{\sc a}
\otimes {\cal C}^4_{\sc b}\otimes {\cal C}^N_{\sc c}$ such that
$r(\la 1_A, 3_B|\rho |1_A, 3_B\ra)=r(\rho)=N$, can be transformed
into the following canonical form by using a reversible local
operation: \be \rho=\sqrt{F}[DC\ \ DB\ \ DA\ \ D \ \ C\ \ B\ \ A\
\ I]^{\dag} [DC\ \ DB\ \ DA\ \ D \ \ C\ \ B\ \ A\ \ I]\sqrt{F},
\ee where $A$, $B$, $C$, $D$, $F$ and the identity $I$ are
$N\times N$ matrices acting on ${\cal C}_{\sc c}^N$ and satisfy
the following relations: $[A,\ A^{\dag}]=[B,\ B^{\dag}]=[C,\
C^{\dag}]=[D,\ D^{\dag}]=[B,\ A]=[B,\ A^{\dag}] =[C,\ A]=[C,\
A^{\dag}]=[D,\ A]=[D,\ A^{\dag}]=[C,\ B]=[C,\ B^{\dag}]=[D,\
B]=[D,\ B^{\dag}]= [D,\ C]=[D,\ C^{\dag}]=0$ and $F=F^{\dag}$
($\dag$ stands for the transposition and conjugate).

{\bf Proof.}\ In the considered basis a density matrix $\rho$ can
be always written as: \be\label{r} \rho=\left(
\begin{array}{cccccccc}
E_1&E_{12}&E_{13}&E_{14}&E_{15}&E_{16}&E_{17}&E_{18}\\
E_{12}^{\dag}&E_2&E_{23}&E_{24}&E_{25}&E_{26}&E_{27}&E_{28}\\
E_{13}^{\dag}&E_{23}^{\dag}&E_{3}&E_{34}&E_{35}&E_{36}&E_{37}&E_{38}\\
E_{14}^{\dag}&E_{24}^{\dag}&E_{34}^{\dag}&E_4&E_{45}&E_{46}&E_{47}&E_{48}\\
E_{15}^{\dag}&E_{25}^{\dag}&E_{35}^{\dag}&E_{45}^{\dag}&E_5&E_{56}&E_{57}&E_{58}\\
E_{16}^{\dag}&E_{26}^{\dag}&E_{36}^{\dag}&E_{46}^{\dag}&E_{56}^{\dag}&E_6&E_{67}&E_{68}\\
E_{17}^{\dag}&E_{27}^{\dag}&E_{37}^{\dag}&E_{47}^{\dag}&E_{57}^{\dag}&E_{67}^{\dag}&E_7&E_{78}\\
E_{18}^{\dag}&E_{28}^{\dag}&E_{38}^{\dag}&E_{48}^{\dag}&E_{58}^{\dag}&E_{68}^{\dag}&E_{78}^{\dag}&E_8\\
\end{array}
\right),
\ee
where $E's$ are $N \times N$ matrices, $r(E_8)=N$.
After the projection $\tilde{\rho}=\la 1_A|\rho|1_A\ra$, we obtain
\be
\tilde{\rho}=\la1_A|\rho|1_A\ra
=\left(
\begin{array}{cccc}
E_5&E_{56}&E_{57}&E_{58}\\
E_{56}^{\dag}&E_6&E_{67}&E_{68}\\
E_{57}^{\dag}&E_{67}^{\dag}&E_7&E_{78}\\
E_{58}^{\dag}&E_{68}^{\dag}&E_{78}^{\dag}&E_8
\end{array}
\right).
\ee

$\tilde{\rho}$ is now a state in ${\cal C}^4_{\sc b} \otimes {\cal
C}^N_{\sc c}$ with $r(\tilde{\rho})=r(\rho)=N$. As every principal
minor determinant of $\tilde \rho^{t_B}$ ($\tilde \rho^{t_C}$) is
some principal minor determinant of $\rho$, the fact that $\rho$
is PPT implies that $\tilde \rho$ is also PPT, $\tilde{\rho}\ge
0$. After performing a reversible local non-unitary ``filtering"
$\frac{1}{\sqrt{E_8}}$ on ${\cal C}^N_{\sc c}$ and using the Lemma
5 in \cite{hlpre} we can express the matrix $\tilde{\rho}$ as
\be
\tilde{\rho}=\left(
\begin{array}{cccc}
C^{\dag}C&C^{\dag}B&C^{\dag}A&C^{\dag}\\
B^{\dag}C&B^{\dag}B&B^{\dag}A&B^{\dag}\\
A^{\dag}C&A^{\dag}B&A^{\dag}A&A^{\dag}\\ C&B&A&I
\end{array}
\right),
\ee
where
$[A,A^{\dag}]=[B,B^{\dag}]=[C,C^{\dag}]=[B,A]=[C,A]=[C,B]=[B,A^{\dag}]=[C,A^{\dag}]=[C,B^{\dag}]=0$.

Similarly, if we consider the projection $\la 3_B|\rho|3_B\ra$ ,
for the same reasons as above we conclude that the resulting
matrix
$$ \ba{rcl}
\bar{\rho}&=&\la 3_B|\rho| 3_B\ra\\[5mm]
&=&\left(
\begin{array}{cc}
E_4&E_{48}\\
E_{48}^{\dag}&E_8
\end{array}
\right)
=\left(
\begin{array}{cc}
D^{\dag}D&D^{\dag}\\D&I
\end{array}
\right)
\ea
$$
where $[D,D^{\dag}]=0$.

Hence, after performing the local filtering operation
$\frac{1}{\sqrt{E_8}}$, the matrix $\rho$ now has the form:
\be\label{rm}
\rho=\left(
\begin{array}{cccccccc}
E_1&E_{12}&E_{13}&E_{14}&E_{15}&E_{16}&E_{17}&E_{18}\\
E_{12}^{\dag}&E_2&E_{23}&E_{24}
&E_{25}&E_{26}&E_{27}&E_{28}\\
E_{13}^{\dag}&E_{23}^{\dag}&E_3&E_{34}
&E_{35}&E_{36}&E_{37}&E_{38}\\
E_{14}^{\dag}&E_{24}^{\dag}&E_{34}^{\dag}&D^{\dag}D&E_{45}
&E_{46}&E_{47}&D^{\dag}\\
E_{15}^{\dag}&E_{25}^{\dag}&E_{35}^{\dag}&E_{45}^{\dag}&C^{\dag}C
&C^{\dag}B&C^{\dag}A&C^{\dag}\\
E_{16}^{\dag}&E_{26}^{\dag}&E_{36}^{\dag}&E_{46}^{\dag}&
B^{\dag}C&B^{\dag}B&B^{\dag}A&B^{\dag}\\
E_{17}^{\dag}&E_{27}^{\dag}&E_{37}^{\dag}&E_{47}^{\dag}&
A^{\dag}C&A^{\dag}B&A^{\dag}A&A^{\dag}\\
E_{18}^{\dag}&E_{28}^{\dag}&E_{38}^{\dag}&D&C&B&A&I\\
\end{array}
\right).
\ee
Owing to that $\rho\geq 0$, a vector $|v\ra$ in
${\cal C}^2_{\sc A}\otimes{\cal C}^M_{\sc B}\otimes{\cal C}^N_{\sc C}$
satisfying $\la v|\rho|v\ra=0$ is in the kernel of $\rho$.
It is directly verified that the following vectors are in the kernel:
\be
\ba{ll}
|10\ra|f\ra-|13\ra
C|f\ra,&~~
|11\ra|g\ra-|13\ra B|g\ra,\\[3mm]
|12\ra|h\ra-|13\ra A|h\ra,&~~ |03\ra|k\ra-|13\ra D|k\ra,
\ea
\ee
for all $|f\ra,~|g\ra,~|h\ra,~|k\ra\in {\cal C}^N_{\sc C}$. This implies
\be\label{imp}
\ba{lll} E_{45}=D^{\dag}C,&~~E_{46}=D^{\dag}B,&~~
E_{47}=D^{\dag}A,\\[3mm]
E_{15}=E_{18}C,&~~
E_{25}=E_{28}C,&~~E_{35}=E_{38}C,\\[3mm]
E_{16}=E_{18}B,&~~ E_{26}=E_{28}B,&~~ E_{36}=E_{38}B,\\[3mm]
E_{17}=E_{18}A,&~~ E_{27}=E_{28}A,&~~ E_{37}=E_{38}A,\\[3mm]
E_{14}=E_{18}D,&~~ E_{24}=E_{28}D,&~~ E_{34}=E_{38}D.
\ea \ee

Substituting (\ref{imp}) into (\ref{rm}) and taking a partial
transposition of $\rho$ with respect to the sub-system ${\sc a}$,
we have \be \rho^{t_{\sc a}}=\left(
\begin{array}{cccccccc}
E_1&E_{12}&E_{13}&E_{18}D&C^{\dag}E_{18}^{\dag}
&C^{\dag}E_{28}^{\dag}&C^{\dag}E_{38}^{\dag}&C^{\dag}D\\
E_{12}^{\dag}&E_2&E_{23}&E_{28}D&
B^{\dag}E_{18}^{\dag}&B^{\dag}E_{28}^{\dag}&B^{\dag}E_{38}^{\dag}&B^{\dag}D\\
E_{13}^{\dag}&E_{23}^{\dag}&E_3&
E_{38}D&A^{\dag}E_{18}^{\dag}&A^{\dag}E_{28}^{\dag}&A^{\dag}E_{38}^{\dag}&A^{\dag}D\\
D^{\dag}E_{18}^{\dag}&D^{\dag}E_{28}^{\dag}&D^{\dag}E_{38}^{\dag}&D^{\dag}D&E_{18}^{\dag}
&E_{28}^{\dag}&E_{38}^{\dag}&D\\
E_{18}C&E_{18}B&E_{18}A&E_{18}&C^{\dag}C
&C^{\dag}B&C^{\dag}A&C^{\dag}\\
E_{28}C&E_{28}B&E_{28}A&E_{28}
&B^{\dag}C&B^{\dag}B&B^{\dag}A&B^{\dag}\\
E_{38}C&E_{38}B&E_{38}A&E_{38}
&A^{\dag}C&A^{\dag}B&A^{\dag}A&A^{\dag}\\
D^{\dag}C&D^{\dag}B&D^{\dag}A&D^{\dag}&C&B&A&I
\end{array}
\right).
\ee

Since the partial transposition with respect to the sub-system
${\sc a}$ is positive, $\rho^{t_{\sc a}}\ge 0$, and it does not
change $\la 1_{\sc a}|\rho|1_{\sc a}\ra$, we still have
$|10\ra|f\ra-|13\ra C|f\ra\in k(\rho^{t_{\sc a}})$. This gives
rise to the following equalities:
$E_{18}^{\dag}=DC,~E_{28}^{\dag}=DB,~E_{38}^{\dag}=DA$,
$E_{18}=C^{\dag}D^{\dag},~E_{28}=B^{\dag}D^{\dag},~E_{38}=A^{\dag}D^{\dag}$.
$\rho$ is then of the following form:
$$
\ba{l}
\left(
\begin{array}{cccccccc}
E_1\!&\!E_{12}\!&\!E_{13}\!&\!
C^{\dag}D^{\dag}D\!&\!C^{\dag}D^{\dag}C
\!&\!C^{\dag}D^{\dag}B\!&\!C^{\dag}D^{\dag}A
\!&\!C^{\dag}D^{\dag}\\
E_{12}^{\dag}&E_2&E_{23}&B^{\dag}D^{\dag}D\!&\!B^{\dag}D^{\dag}
C\!&\!B^{\dag}D^{\dag}B\!&\!B^{\dag}D^{\dag}A
\!&\!B^{\dag}D^{\dag}\\
E_{13}^{\dag}\!&\!E_{23}^{\dag}\!&\!E_3\!&\!
A^{\dag}D^{\dag}D\!&\!A^{\dag}D^{\dag}C\!&\!A^{\dag}D^{\dag}B\!&\!A^{\dag}D^{\dag}A
\!&\!A^{\dag}D^{\dag}\\
D^{\dag}DC\!&\!D^{\dag}DB\!&\!D^{\dag}DA\!&\!
D^{\dag}D\!&\!D^{\dag}C\!&\!D^{\dag}B\!&\!D^{\dag}A\!&\!D^{\dag}\\
C^{\dag}DC\!&\!C^{\dag}DB\!&\!C^{\dag}DA\!&\!
C^{\dag}D\!&\!C^{\dag}C\!&\!C^{\dag}B\!&\!C^{\dag}A
\!&\!C^{\dag}\\
B^{\dag}DC\!&\!B^{\dag}DB\!&\!B^{\dag}DA\!&\!
B^{\dag}D\!&\!B^{\dag}C\!&\!B^{\dag}B\!&\!B^{\dag}A\!&\!B^{\dag}\\
A^{\dag}DC\!&\!A^{\dag}DB\!&\!A^{\dag}DA\!&\!
A^{\dag}D\!&\!A^{\dag}C\!&\!A^{\dag}B\!&\!A^{\dag}A\!&\!A^{\dag}\\
DC\!&\!DB\!&\!DA\!&\!D\!&\!C\!&\!B\!&\!A\!&\!I
\end{array}
\right).\\[20mm]
\ea
$$
Set
$$
X=\left(
\begin{array}{cccccc}
E_{13}&C^{\dag}D^{\dag}D&C^{\dag}D^{\dag}C&C^{\dag}D^{\dag}B&C^{\dag}D^{\dag}A&C^{\dag}D^{\dag}\\
E_{23}&B^{\dag}D^{\dag}D&B^{\dag}D^{\dag}C&B^{\dag}D^{\dag}B&B^{\dag}D^{\dag}A&B^{\dag}D^{\dag}
\end{array}
\right),
$$
$$
Y=\left(
\begin{array}{cc}
E_1&E_{12}\\
E_{12}^{\dag}&E_2
\end{array}
\right),
$$
and
$$
\rho_6=\Sigma+{\rm diag}(\Delta,\ 0,\ 0,\ 0,\ 0,\ 0),
$$ \\
where $$ \Sigma=\left(
\begin{array}{cccccc}
A^{\dag}D^{\dag}DA&A^{\dag}D^{\dag}D&A^{\dag}
D^{\dag}C&A^{\dag}D^{\dag}B&A^{\dag}D^{\dag}A&A^{\dag}D^{\dag}\\
D^{\dag}DA&D^{\dag}D&D^{\dag}C&D^{\dag}B&D^{\dag}A&D^{\dag}\\
C^{\dag}DA&C^{\dag}D&C^{\dag}C&C^{\dag}B&C^{\dag}A&C^{\dag}\\
B^{\dag}DA&B^{\dag}D&B^{\dag}C&B^{\dag}B&B^{\dag}A&B^{\dag}\\
A^{\dag}DA&A^{\dag}D&A^{\dag}C&A^{\dag}B&A^{\dag}A&A^{\dag}\\
DA&D&C&B&A&I
\end{array}
\right),
$$
\be\label{delta} \Delta=E_3-A^{\dag}D^{\dag}DA, \ee
diag$(A_1,A_2,...,A_m)$ denotes a diagonal block matrix with
blocks $A_1,A_2,...,A_m$. $\rho$ can then be written in the
following partitioned matrix form:
\[
\rho=\left(
\begin{array}{cc}
Y&X\\
X^{\dag}&\rho_6
\end{array}
\right).
\]
As $\Sigma$ possesses the following  5N kernel vectors: \be
\ba{ll} (\la f|, \ 0,\ 0,\ 0,\ 0, -\la f|A^\dag D^\dag  )^T,\ \
&~~~
(0, \la g|,0,\ 0,\ 0, -\la g|D^\dag )^T,\\[3mm]
(0,\ 0,\la h|, \ 0,\ 0,\ -\la h|C^\dag )^T,\ \ &~~~
(0,\ 0,\ 0,\la i|,\ 0,\ -\la i|B^\dag )^T,\\[3mm]
\ \ (0,\ 0,\ 0,\ 0,\la j|,\ -\la j|A^\dag )^T, \ea \ee for
arbitrary $|f\ra,\ |g\ra,\ |h\ra,\ |i\ra,\ |j\ra \in{\cal C}^N_{\sc C}$,
so the kernel $K(\Sigma)$ has at least dimension $5N$. On
the other hand $r(\Sigma)+k(\Sigma)=6N$, therefore $r(\Sigma)\leq
N$. While the range of $\Sigma$ has at least dimension $N$ due to
the identity entry on the diagonal. Hence we have $r(\Sigma)=N$.

Taking into account that $r(\rho_6)\leq r(\rho)=N$, it is easy to see that
$r(\rho_6)=N$. As the rank of $\rho_6$ is $N$, by making the following
elementary row transformations on the matrix $\rho_6$,
\be\label{add} \left(
\begin{array}{cccccc}
I&0&0&0&0&-A^{\dag}D^{\dag}\\
0&I&0&0&0&-D^{\dag}\\
0&0&I&0&0&-C^{\dag}\\
0&0&0&I&0&-B^{\dag}\\
0&0&0&0&I&-A^{\dag}\\
0&0&0&0&0&I
\end{array}
\right)\rho_6= \left(
\begin{array}{cccccc}
\Delta&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
DA&D&C&B&A&I
\end{array}
\right). \ee we deduce that $\Delta=0$, i.e.,
$E_3=A^{\dag}D^{\dag}DA$.

Similarly, $|02\ra |f\ra -|13\ra DA|f\ra$
is also in the kernel for all $|f\ra\in{\cal C}^N_{\sc C}$,
from which we have $E_{13}=C^{\dag}D^{\dag}DA$,
$E_{23}=B^{\dag}D^{\dag}DA$ and the matrix $\rho$ becomes
$$
\left(
\begin{array}{cccccccc}
E_1\!&\!E_{12}\!&\! C^{\dag}D^{\dag}DA\!&\!
C^{\dag}D^{\dag}D\!&\!C^{\dag}D^{\dag}C
\!&\!C^{\dag}D^{\dag}B\!&\!C^{\dag}D^{\dag}A
\!&\!C^{\dag}D^{\dag}\\
E_{12}^{\dag}&E_2&B^{\dag}D^{\dag}DA&B^{\dag}D^{\dag}D\!&\!
B^{\dag}D^{\dag}C\!&\!B^{\dag}D^{\dag}B\!&\!B^{\dag}D^{\dag}A
\!&\!B^{\dag}D^{\dag}\\
A^{\dag}D^{\dag}DC\!&\!A^{\dag}D^{\dag}DB\!&\!A^{\dag}D^{\dag}DA\!&\!
A^{\dag}D^{\dag}D\!&\!A^{\dag}D^{\dag}C\!&\!A^{\dag}D^{\dag}B\!&\!A^{\dag}D^{\dag}A
\!&\!A^{\dag}D^{\dag}\\
D^{\dag}DC\!&\!D^{\dag}DB\!&\!D^{\dag}DA\!&\!
D^{\dag}D\!&\!D^{\dag}C\!&\!D^{\dag}B\!&\!D^{\dag}A\!&\!D^{\dag}\\
C^{\dag}DC\!&\!C^{\dag}DB\!&\!C^{\dag}DA\!&\!
C^{\dag}D\!&\!C^{\dag}C\!&\!C^{\dag}B\!&\!C^{\dag}A
\!&\!C^{\dag}\\
B^{\dag}DC\!&\!B^{\dag}DB\!&\!B^{\dag}DA\!&\!
B^{\dag}D\!&\!B^{\dag}C\!&\!B^{\dag}B\!&\!B^{\dag}A\!&\!B^{\dag}\\
A^{\dag}DC\!&\!A^{\dag}DB\!&\!A^{\dag}DA\!&\!
A^{\dag}D\!&\!A^{\dag}C\!&\!A^{\dag}B\!&\!A^{\dag}A\!&\!A^{\dag}\\
DC\!&\!DB\!&\!DA\!&\!D\!&\!C\!&\!B\!&\!A\!&\!I
\end{array}
\right).\\
$$
Similarly, using the same method  as for $E_3$ and $E_{23}$
respectively, we can derive $E_2=B^{\dag}D^{\dag}DB$,
$E_{12}=C^{\dag}D^{\dag}DB$, $E_1=C^{\dag}D^{\dag}DC$. $\rho$ then
is the following form:
$$
\ba {l} \left(
\begin{array}{cccccccc}
C^{\dag}D^{\dag}DC\!&\!C^{\dag}D^{\dag}DB\!&\!C^{\dag}D^{\dag}DA\!&\!
C^{\dag}D^{\dag}D\!&\!C^{\dag}D^{\dag}C
\!&\!C^{\dag}D^{\dag}B\!&\!C^{\dag}D^{\dag}A
\!&\!C^{\dag}D^{\dag}\\
B^{\dag}D^{\dag}DC&B^{\dag}D^{\dag}DB&B^{\dag}D^{\dag}DA&B^{\dag}D^{\dag}
D\!&\!B^{\dag}D^{\dag}C\!&\!B^{\dag}D^{\dag}B\!&\!B^{\dag}D^{\dag}A
\!&\!B^{\dag}D^{\dag}\\
A^{\dag}D^{\dag}DC\!&\!A^{\dag}D^{\dag}DB\!&\!A^{\dag}D^{\dag}DA\!&\!
A^{\dag}D^{\dag}D\!&\!A^{\dag}D^{\dag}C\!&\!A^{\dag}D^{\dag}B\!&\!A^{\dag}D^{\dag}A
\!&\!A^{\dag}D^{\dag}\\
D^{\dag}DC\!&\!D^{\dag}DB\!&\!D^{\dag}DA\!&\!
D^{\dag}D\!&\!D^{\dag}C\!&\!D^{\dag}B\!&\!D^{\dag}A\!&\!D^{\dag}\\
C^{\dag}DC\!&\!C^{\dag}DB\!&\!C^{\dag}DA\!&\!
C^{\dag}D\!&\!C^{\dag}C\!&\!C^{\dag}B\!&\!C^{\dag}A
\!&\!C^{\dag}\\
B^{\dag}DC\!&\!B^{\dag}DB\!&\!B^{\dag}DA\!&\!
B^{\dag}D\!&\!B^{\dag}C\!&\!B^{\dag}B\!&\!B^{\dag}A\!&\!B^{\dag}\\
A^{\dag}DC\!&\!A^{\dag}DB\!&\!A^{\dag}DA\!&\!
A^{\dag}D\!&\!A^{\dag}C\!&\!A^{\dag}B\!&\!A^{\dag}A\!&\!A^{\dag}\\
DC\!&\!DB\!&\!DA\!&\!D\!&\!C\!&\!B\!&\!A\!&\!I
\end{array}
\right)\\[20mm]
=[\begin{array}{cccccccc} DC& DB&DA&D&C&B&A&I\end{array}]^{\dag}
[DC\,\,\, DB\,\,\, DA\,\,\, D\,\,\, C\,\,\, B\,\,\, A\,\,\, I].\\[3mm]
\ea
$$

The commutative relations $[A,\ D]=[B,\ D]=[C,\ D]=[A,\
D^{\dag}]=[B,\ D^{\dag}]=[C,\ D^{\dag}]=0$ follow from the
positivity of all partial transpositions of $\rho$. We first
consider: \be \rho^{t_{\sc b}}=\left(
\begin{array}{cccccccc}
C^{\dag}D^{\dag}DC\!&\!B^{\dag}D^{\dag}DC\!&\!A^{\dag}D^{\dag}DC\!&\!
D^{\dag}DC\!&\!C^{\dag}D^{\dag}C
\!&\!B^{\dag}D^{\dag}C\!&\!A^{\dag}D^{\dag}C
\!&\!D^{\dag}C\\
C^{\dag}D^{\dag}DB\!&\!B^{\dag}D^{\dag}DB\!&\!A^{\dag}D^{\dag}DB\!&\!
D^{\dag}DB\!&\!C^{\dag}D^{\dag}B\!&\!B^{\dag}D^{\dag}B\!&\!A^{\dag}D^{\dag}B
\!&\!D^{\dag}B\\
C^{\dag}D^{\dag}DA\!&\!B^{\dag}D^{\dag}DA\!&\!A^{\dag}D^{\dag}DA\!&\!
D^{\dag}DA\!&\!C^{\dag}D^{\dag}A\!&\!B^{\dag}D^{\dag}A\!&\!A^{\dag}D^{\dag}A
\!&\!D^{\dag}A\\
C^{\dag}D^{\dag}D\!&\!B^{\dag}D^{\dag}D\!&\!A^{\dag}D^{\dag}D\!&\!
D^{\dag}D\!&\!C^{\dag}D^{\dag}\!&\!B^{\dag}D^{\dag}\!&\!A^{\dag}D^{\dag}\!&\!D^{\dag}\\
C^{\dag}DC\!&\!B^{\dag}DC\!&\!A^{\dag}DC\!&\!
DC\!&\!C^{\dag}C\!&\!B^{\dag}C\!&\!A^{\dag}C
\!&\!C\\
C^{\dag}DB\!&\!B^{\dag}DB\!&\!A^{\dag}DB\!&\!
DB\!&\!C^{\dag}B\!&\!B^{\dag}B\!&\!A^{\dag}B\!&\!B\\
C^{\dag}DA\!&\!B^{\dag}DA\!&\!A^{\dag}DA\!&\!
DA\!&\!C^{\dag}A\!&\!B^{\dag}A\!&\!A^{\dag}A\!&\!A\\
C^{\dag}D\!&\!B^{\dag}D\!&\!A^{\dag}D\!&\!D\!&\!C^{\dag}\!&\!B^{\dag}\!&\!A^{\dag}\!&\!I
\end{array}
\right).
\ee
Due to the positivity, the matrix $\rho^{t_B}$ must
possess the kernel vector $|03\ra |h\ra -|13\ra D |h\ra$, which
implies that $[A,\ D]=[B,\ D]=[C,\ D]=0$.The matrix $\rho^{t_B}$
can be then written as:
\[
\rho^{t_B}=\left(
\begin{array}{c}
D^{\dag}C\\
D^{\dag}B\\
D^{\dag}A\\
D^{\dag}\\
C\\
B\\
A\\
I
\end{array}
\right) \left(
\begin{array}{cccccccc}
C^{\dag}D&B^{\dag}D&A^{\dag}D&D&C^{\dag}&B^{\dag}&A^{\dag}&I
\end{array}
\right),
\]
which implies automatically the positivity.

From the positivity of $\rho^{t_{AB}}$,
 \[
\rho^{t_{AB}}=\left(
\begin{array}{cccccccc}
C^{\dag}D^{\dag}DC&B^{\dag}D^{\dag}DC&A^{\dag}D^{\dag}DC&D^{\dag}DC&C^{\dag}DC&B^{\dag}DC&A^{\dag}DC&DC\\
C^{\dag}D^{\dag}DB&B^{\dag}D^{\dag}DB&A^{\dag}D^{\dag}DB&D^{\dag}DB&C^{\dag}DB&B^{\dag}DB&A^{\dag}DB&DB\\
C^{\dag}D^{\dag}DA&B^{\dag}D^{\dag}DA&A^{\dag}D^{\dag}DA&D^{\dag}DA&C^{\dag}DA&B^{\dag}DA&A^{\dag}DA&DA\\
C^{\dag}D^{\dag}D&B^{\dag}D^{\dag}D&A^{\dag}D^{\dag}D&D^{\dag}D&C^{\dag}D&B^{\dag}D&A^{\dag}D&D\\
C^{\dag}D^{\dag}C&B^{\dag}D^{\dag}C&A^{\dag}D^{\dag}C&D^{\dag}C&C^{\dag}C&B^{\dag}C&A^{\dag}C&C\\
C^{\dag}D^{\dag}B&B^{\dag}D^{\dag}B&A^{\dag}D^{\dag}B&D^{\dag}B&C^{\dag}B&B^{\dag}B&A^{\dag}B&B\\
C^{\dag}D^{\dag}A&B^{\dag}D^{\dag}A&A^{\dag}D^{\dag}A&D^{\dag}A&C^{\dag}A&B^{\dag}A&A^{\dag}A&A\\
 C^{\dag}D^{\dag}&B^{\dag}D^{\dag}&A^{\dag}D^{\dag}&D^{\dag}&C^{\dag}&B^{\dag}&A^{\dag}&I
\end{array}
\right),
\]
we have that $|03\ra |h\ra -|13\ra D^{\dag} |h\ra$ is a kernel
vector, which results in $[A,\ D^{\dag}]=[B,\ D^{\dag}]=[C,\
D^{\dag}]=0$. $\rho^{t_{AB}}$ is then of the form:
\[
\rho^{t_{AB}}=\left(
\begin{array}{c}
DC\\
DB\\
DA\\
D\\
C\\
B\\
A\\
I
\end{array}
\right) \left(
\begin{array}{cccccccc}
C^{\dag}D^{\dag}&B^{\dag}D^{\dag}&A^{\dag}D^{\dag}&D^{\dag}&C^{\dag}&B^{\dag}&A^{\dag}&I
\end{array}
\right).
\]
This form assures positive definiteness, and concludes the proof
of the Lemma. $\Box$

Using Lemma 1 we can prove the following Theorem:

{\bf Theorem 1.}\ \ A PPT-state $\rho$ in ${\cal C}^2_{\sc A}
\otimes {\cal C}^4_{\sc B} \otimes {\cal C}^N_{\sc C}$ with
$r(\rho)=N$ is separable if there exists a product vector $|e_A,\
f_B\ra$ of ${\cal C}^2_{\sc A} \otimes {\cal C}^4_{\sc B}$ such
that $r(\la e_A,\ f_B|\rho |e_A,\ f_B\ra)=N$.

{\bf Proof.}\ \ According to the Lemma 1 the PPT state $\rho$ can
be written as
\[
\rho=
\left(
\begin{array}{c}
C^{\dag}D^{\dag}\\B^{\dag}D^{\dag}\\A^{\dag}D^{\dag}\\
D^{\dag}\\C^{\dag}\\B^{\dag}\\A^{\dag}\\I
\end{array}
\right)
\left(
\begin{array}{cccccccc}
DC&DB&DA&D&C&B&A&I
\end{array}
\right).
\]
Since all $A$, $A^\dag$, $B$, $B^\dag$, $C$, $C^\dag$, $D$ and
$D^\dag$ commute, they have common eigenvectors $|f_n\ra$. Let
$a_n$, $b_n$, $c_n$ and $d_n$ be the corresponding eigenvalues of
$A$, $B$, $C$ and $D$ respectively. We have
\[ \la f_n|\rho
|f_n\ra= \left(
\begin{array}{c}
c_n^*d_n^*\\[2mm]b_n^*d_n^*\\[2mm]a_n^*d_n^*\\[2mm]d_n^*\\[2mm]c_n^*\\[2mm]
b_n^*\\[2mm]a_n^*\\[2mm]1
\end{array}
\right)
\left(
\begin{array}{cccccccc}
d_nc_n&d_nb_n&d_na_n&d_n&c_n&b_n&a_n&1
\end{array}
\right)\hskip5cm
\]
\[
=\left[\left(\begin{array}{c}d_n^*\\1\end{array}\right)\otimes
\left(\begin{array}{c}c_n^*\\b_n^*\\a_n^*\\1\end{array}\right)\right]
(d_n\,\,\, 1)\otimes(c_n,b_n,a_n,\,1)
=|e_{\sc a},f_{\sc b}\ra\la
e_{\sc a},f_{\sc b}|.
\]
We can thus write $\rho$ as
$$\rho=\sum_{n=1}^N|\psi_n\ra\la \psi_n|\otimes |\phi_n\ra\la \phi_n|
\otimes |f_n\ra\la f_n|,
$$
where
$$
|\psi_n\ra=\left(\begin{array}{c}d_n^*\\ 1\end{array}\right),~~~
|\phi_n\ra=\left(\begin{array}{c}c_n^*\\b_n^*\\a_n^*\\
 1\end{array}\right).
$$
Since the local transformations are reversible, we can apply the
inverse transformations and obtain a decomposition of the initial
state $\rho$ in a sum of projectors onto the product vectors.
Therefore $\rho$ is separable. \hfill $\Box$

The above approach can be extended to the higher dimensional case
${\cal C}^2_{\sc a} \otimes {\cal C}^M_{\sc b}\otimes {\cal
C}^N_{\sc c}$. Similar to Lemma 1, it is straightforward to prove
the following conclusion:

 {\bf Lemma 2. }\ \  Every PPT state $\rho$ in
${\cal C}^2_{\sc a} \otimes {\cal C}^M_{\sc b}\otimes {\cal
C}^N_{\sc c}$ such that $r(\la 1_A, M-1_B|\rho |1_A,
M-1_B\ra)=r(\rho)=N$, can be transformed into the following
canonical form by using a reversible local operation:
$$
\rho=\sqrt{F}[DA_{M-1}\ \cdots\ \ DA_1\ \ D \ \ A_{M-1}\ \cdots\ \
A_1\ \ I]^{\dag} [DA_{M-1}\ \cdots\ \ DA_1\ \ D \ \ A_{M-1}\
\cdots\ \ A_1\ \ I]\sqrt{F},
$$
where $A_1$, $\cdots$, $A_{M-1}$, $D$, $F$ and the identity $I$
are $N\times N$ matrices acting on ${\cal C}_{\sc c}^N$ and
satisfy the following relations: $[A_i,\ {A_j}^{\dag}]=[A_i,\
A_j]=[D,\ D^{\dag}]=[D,\ A_i]=[D,\ {A_i}^{\dag}]=0$,
$i,j=1,2,\cdots,M-1$ and $F=F^{\dag}$ . Set $T=(D\ \
I)\otimes(A_{M-1}\ \ A_{M-2}\ \cdots\ \ A_1\ \ I)$, we have \be
\rho=\sqrt{F} T^{\dag}T \sqrt{F}\ee

Extending Theorem 1 to higher dimensional cases
${\cal C}^2_{\sc a} \otimes {\cal C}^M_{\sc
b}\otimes {\cal C}^N_{\sc c}$, we have:

{\bf Theorem 2.}\ \ A PPT-state $\rho$ in ${\cal C}^2 \otimes
{\cal C}^M \otimes {\cal C}^N$ with $r(\rho)=N$ is separable if
there exists a product vector $|e_A,\ f_B\ra$ such that $r(\la
e_A,\ f_B|\rho |e_A,\ f_B\ra)=N$.

A separable state is always PPT. But a PPT state in ${\cal C}^2
\otimes {\cal C}^M \otimes {\cal C}^N$ is not generally separable.
Our canonical form of PPT states in ${\cal C}^2 \otimes {\cal C}^M
\otimes {\cal C}^N$ composite quantum systems provides a way to
investigate the separability of such states and the structures of
bound entangled states, as all entangled PPT states are bound
entangled states.

\end{document}